OpenStudy (anonymous):

-2x+4y=-5 Write two equations in STANDARD FORM that are equivalent to the given equation

OpenStudy (agreene):

Basically this is asking you to solve for an x=STUFF and y=STUFF pair of equations. Try and do that, let me know if you get stuck on something.

OpenStudy (anonymous):

Just show me how to do this one there are others

OpenStudy (agreene):

-2x+4y=-5 Solving for y 4y=-5+2x y= -5/4 + 1/2x -2x+4y=-5 solving for x -2x=-5-4y x=5/2 + 2y

OpenStudy (anonymous):

Do you understand its in standard form Ax+By=C

OpenStudy (agreene):

That depends on your definition of standard.

OpenStudy (anonymous):

Ax+By=C

OpenStudy (agreene):

But, if you must insist that this is standard. then multiply everything by a constant coefficient and you have your answer.

OpenStudy (anonymous):

Your not solving for x your not solving for y your trying two write two equations In AX+BY==C form not some other form

OpenStudy (agreene):

As I just said, you can multiply the entire equation by any number you choose and get an congruent set.

OpenStudy (anonymous):

Ya ok i see thanks

OpenStudy (agreene):

For instance, if you choose pi as your constant. Ax+By=C is congruent to Apix+Bpiy=Cpi